OpenStudy (anonymous):

use substitution to solve the system: 15x- = 5y + 5 y = 2x + 1 A. (5, 3) B. (5, -2) C (2, 3) D (2, 5)

OpenStudy (anonymous):

@triciaal @ashleydavidson939 @salparadise64

OpenStudy (triciaal):

substitution 15x = 5(2x +1) + 5 15x = 10x + 5 + 5 5x = 10 x = 2

OpenStudy (triciaal):

now you have x = 2 put this in the eqn for y y = 2x + 1 = 2*2 + 1 = 5 solution x =2 y = 5 d =(2,5)